Incident Report

On this page you will find the latest information. This page is also available in Dutch.

[28-09-2017] SOLVED: J-drive is accessible


M1709 5430


SOLVED: J-drive is not accessible

There has been a disruption to the J-drive, which has already been solved. The J-drive is fully operational again.




Begin time:

11:15 hrs.

End time:

10:45 hrs.

[22-08-2017] J-drive temporary inaccessible - SOLVED


M1708 2790


J-drive operational again

There has been a disruption to the J-drive, which has already been solved. The J-drive is fully operational again.



Begin time:


End time:


[17-08-2017] uMail, uWebguide and uTools not available - SOLVED


M1708 2228


Services restored


uMail, uWebguide and uTools are available again.




Begin time:

16.45 hrs

End time:

 20.05 hrs

[28-06-2017] Update: Services largely recovered


M1706 2795


Update: services largely recovered

The disruption in our service that arose last Monday is now largely recovered. We are still working on the search function on the university websites. We expect the search function to be fully functional again tomorrow morning June 29.
Our analysis of the root cause as far as we have made it, is that the network disruption was rooted in the combination of the way DNS, redundancy and authentication were working together. Because of this, the network infrastructure did not react as expected when a disruption occurred. A more detailed analysis of this root cause will be made as part of the Problem-process. The recommendations for improvements that will result from this will be implemented in the near future. 



Begin time:

09:20 hrs


28-06-2017 – 15:00hrs

[09-05-2017] Network disruption

Incident: M1705 1184

Services restored

Network disruption

The malfunction of the switches at the KOG-building has been resolved. The disruption of other services that was a consequence of this have also been resolved. All test of our services have been successful. The ISSC will now look into the cause of this incident.


Unfortunately, we have not been able to provide updates through the ISSC website, as the CMS was unavailable to us. Simliarly, we were unable to provide updates via SMS.

Date: 09-05-2017
Start time: 13:00hrs
End time: 13:30hrs

[24-04-2017] Service storage P-disk disrupted

Incident: M1704 2980

Service storage P-disk restored

P-disk unavailable for staff with user names starting with a-d 

A part of the P-disk storage is currently unavailable. Only staff with user names starting with letter a-d are affected. This may cause an error report during the login process.  In other cases, the files on your P-disk may be invisible in the file explorer. We are working hard to restore the service.

Update 10:15: A new server is being installed to replace the disturbed part of the p-disks. 

Update 11:15: The new server is up and running. We expect that at approximately 13:00 all staff can view and use their files on the p-disk again.

Update 13:05: Unfortunately, configuring the server is taking more time then expected. We will have an update as soon as everything is up and running and tested properly. 
Update 14:35: The server is up and running, we are testing with users who reported this issue whether everything is working properly again.

Update 16:00: The service has been restored. After a reboot the files on your P:-drive should be available again. The backup and the previous versions of your files have not been affected by the disruption of the service.

Date: 24-04-2017
Start time:

24-04 23:00 h

End time: 25-04 16:00

[03-04-2017] Login issues with uMail

Incident: M1704 0235

Major incident 'Login issues with uMail'
>> services restored

Login issues with uMail

At this moment it’s possible to login to uMail again.

Further applications such as uTools/uWeb/KMS also function correctly.

Date: 03-04-2017
Begin time:

15:16 hrs

End time: 18:40 hrs

[03-04-2017] Network login problems

Incident: M1704 0022

Major incident 'Network login problems'
>> services restored

Network login problems

Due to an interruption last weekend, some computers have lost network connectivity.

Rebooting the computer should offer a solution. Sorry for the inconvenience.

Date: 03-04-2017
Begin time:

09:00 hrs

End time: 13:06 hrs

[27-03-2017] DNS disruption

Incident: M1703 3814

Major incident 'DNS disruption'
>> services restored

Disturbance DNS

Because of disruption on the DNS server, different services were unavailable.

The disruption had impact on the following services: Radius server, DNS server, mail sever, Planon, network storage, website.


Latest update (15.45 hrs): 
All services are fully restored and operational again.

Date: 27-03-2017
Start time:

14:35 hrs

End time: 15:45 hrs

[08-02-2017] Web Lectures service restored

Incident: M1702 1262
Status: Service restored
Date: 08-02-2017
Start time:


End time: 16:30

[26-01-2017] Major incident 'J:\-drive limited approachable'


M1701 3866

Status: Major incident 'J:\-drive limited approachable'

>> Service restored


J:\-drive limited approachable

This morning, there were storage problems. Causing portions of the J:\-drive, to be poor or even not reachable.

Latest update (13:50 hrs):

Reference numbers incidents: M1611 0509, M1612 1394, M1701 1596, M1611 2588, M1701 3866

Reference numbers problem: P1611 0002


Following the last update about the storage issues, we provide you with a new update:


Early November the ISSC identified the first delays observed in the storage (J:\drive).

After this, there has been several (smaller) noticable disruptions of service. From the beginning we were in contact with our supplier Dell to search for the solution. Within the ISSC there is a group of employees with various fields of knowledge that came together on a daily basis in order to solve the problem. The last weeks, analysis resulted in a series of actions from Dell in order to improve the storage environment. These took place in the evenings and weekends and were carried out with as little impact as possible to users. The work performed had to do with changes to the software / firmware of the storage components, but also to the network (iSCSI traffic). Today the storage was available again after 10 minutes of disruption.


The disruption of January 26, 2017, showed that recent changes haven’t produced the desired results. The ISSC and the supplier are working on solving this with the highest priority.


The major call with reference number: M1701 3866 will be closed and continue in Problem: P1611 0002.


Date: 26-01-2016
Start time:

10:50 hrs

End time: 13:50 hrs

Last Modified: 24-10-2017